OpenStudy (anonymous):

A 0.50 kg ball traveling at 6.0 m/s collides head-on with a 1 kg ball moving in the opposite direction at a velocity of -12 m/s. The 0.50 kg ball moves away at -14 m/s after the collision. FInd the velocity of the second ball. PLEASE EXPLAIN!

OpenStudy (anonymous):

in this case you can use the conservation of linear momentum. before \[m \times V1 +M \times V2=Pb\] after \[m \times V1' +M \times V2'=Pa\] by conservation of linear momentum Pb=Pa

OpenStudy (anonymous):

i am so confused about everything. can u please start from the beginning and plug the numbers in the equation you wrote

OpenStudy (anonymous):

The principle of conservation of momentum says that the momentum of the system must be the same before and after the collision. ok?

OpenStudy (anonymous):

oh thanks! i get it

OpenStudy (anonymous):

wait so does that mean that the second ball's velocity is 6 m/s too!?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what is it then?

OpenStudy (anonymous):

\[m \times V1+M \times V2= m \times V1'+M \times V2'\]

OpenStudy (anonymous):

m=0.5 kg M = 1kg V1=6.0 m/s V2= -12 m/s V1'= -14 m/s V2' unknown

OpenStudy (anonymous):

is the answer -2?